The area of a rectangle Garden is 140 mē the width is 3 m longer than one half of the link find the length and width of the garden
Answer by Boreal(15235)

length=x
width is (x/2)+3
x*(x+6)/2=140, putting (x/2)+3 over a common denominator.
x(x+6)=280
x^2+6x-280=0
(x+20)(x-14)=0
x=14 m length, only positive root
width is 3 more than half or 10 m
14 m x 10 m